# Generalization of Diffusion Models Arises with a Balanced Representation Space

### v1

@@ 3.3  Case 3: Hybrid of Memorization and Generalization with Imbalanced Data
Large-scale diffusion datasets often contain duplicates due to imperfect curation or heterogeneous aggregation [carlini2023extracting]; such samples are more easily memorized [somepalli2023understanding] (See Section˜A.4 for more discussions). We model this by allowing duplicated (rank-1) clusters alongside well-sampled, nondegenerate clusters, so the DAE can admit local minimizers that mix memorization and generalization blocks:
@@ 4.1  Representation-Based Memorization Detection
Building on our theoretical insights, we investigate whether memorization can be detected directly from internal representations. Prior work has largely focused on how certain prompts trigger memorization and often relies on those for detection [wen2024detecting, jeonunderstanding, ren2024unveiling]. Representative approaches include: (i) Density-based: detecting samples that are generated disproportionately frequently under a prompt [carlini2023extracting]; and (ii) Norm-based: comparing conditional vs. unconditional scores [wen2024detecting] and (iii) Attention-based: locating anomaly in the cross-attention induced by memorized prompts [hintersdorf2024finding, chen2024exploringBE]. A notable exception is (iv) a landscape-based method of [ross2024geometric], which evaluates memorization using local score-function geometry around a generated sample. Their method makes detection prompt-free, but is still based on output space.
@@
[carlini2023extracting] 0.498 0.020 3.724 N/A N/A
[wen2024detecting] 0.986 0.961 0.134 N/A N/A
[hintersdorf2024finding] 0.957 0.500 0.009 N/A N/A
[ross2024geometric] $\checkmark$ 0.956 0.915 0.545 0.971 0.528 0.031 0.713 0.013 0.071
@@ 4.2  Representation-Space Steering for Interpretable Image Editing
We note that we do not intend to compete with existing outstanding steering methods, as several such approaches have already demonstrated impressive empirical success [hertz2023prompt, zhang2023adding, gandikota2024concept, kadkhodaie2024feature, chen2024exploring]. Rather, our focus is on showing how steering reveals the dual relationship between representation structure and generation behavior. In particular, when the model generalizes, its representations form compositional and interpretable spaces, enabling continuous and controllable edits.
@@ 5.1  Analysis of Learning Diffusion Models with Specific Model Parameterizations
There has been a large body of work analyzing the learning of diffusion models [wang2024evaluating, chen2023score, liang2025low, yang2024few]. Recently, more attention has turned to when and how they overfit or generalize: [li2023generalization, bonnaire2025diffusion] use random-feature assumptions, while [wang2024diffusion, buchanan2025edgememorizationdiffusionmodels] studied empirical denoisers with learnable attractors; [wu2025taking, chen2025interpolation, ye2025provable] investigated smoothing effects induced by learning rates and weight decay that promote generalization. These works are theoretically rigorous but often lack real-world validation.
@@ 5.2  Memorization and Generalization with Analytical DMs
Recent works characterize how architectural or inductive biases can push empirical scores toward more generalizable solutions. For instance, [scarvelis2023closed, lukoianov2025locality] constructed closed-form diffusion models from data; [niedoba2024nearest, niedoba2024towards, kamb2025an, wangseeds] imposed locality or translation-equivariance constraints to mimic U-Net behavior; and [kadkhodaie2023generalization, an2024inductive, floros2025anisotropy] analyzed architectural biases of CNNs and DiTs. [baptista2025memorization] empirically evaluated the impact of various regularization schemes.
@@
[radhakrishnan2020overparameterized, ambrogioni2024search, pham2025memorization] model imperfect training and sampling jointly as an AM recall process, viewing novel image generation as new attraction basins and memorization as perfect recalls [biroli2024dynamical, lyu2025resolving]. However, this perspective can understate the role of learned neural networks in enabling generalization.
@@ 5.3  Studies on Representation Learning of Diffusion Models
Concurrent work studies co-emerging representation learning [kwon2022diffusion, hanfeature] with distribution learning in diffusion models. As recent works [chen2024deconstructing, xiang2025ddae++] re-emphasize that the diffusion objective is fundamentally a self-supervised autoencoder loss [vincent2010stacked, vincent2011connection, bengio2013generalized], which induces encoder-decoder behavior [chen2024deconstructing] and the model autonomously learns informative features for downstream tasks [baranchuk2021label, xiang2023denoising, zhao2023unleashing]. Moreover, supervising the representations can accelerate training [yu2024representation, wang2025learning, singh2025matters], and different representation behaviors correlate with different degrees of overfitting [li2025understanding].
@@ 6  Discussion
Our Final Thoughts Diffusion models generalize mainly because, under the self-supervised denoising objective, neural networks are driven to learn and exploit the underlying (low-dimensional) structures of the data distribution. This capability is reflected internally through the emergence of semantic low-dimensional representation spaces (or auto-encoding [bengio2013generalized]): the network effectively processes/projects noisy inputs with respect to learned structures, which underlie its compressing and denoising behavior [li2025back, kadkhodaie2023generalization, ulyanov2018deep]. In this sense, learning a good (balanced and semantic) representation is a useful indicator of model generalizability. For our theoretical analysis, we focused on a simplified setting designed for analytical tractability, which nonetheless yields foundational intuition into how real-world models function. Specifically, by examining a two-layer ReLU DAE trained on data drawn from a separable MoG distribution—a setting where denoising, score learning, and representation learning are all well defined—we demonstrate that the model learns to map inputs from the same Gaussian mode to the same ReLU mask (i.e., the same subset of active neurons). This selectivity [balestriero2025geometry, song2025selective] leads a simple, interpretable form of representation learning, and we view it as a fundamental reason why neural networks effectively adapt to structured data distributions.
@@ Acknowledgments
We acknowledge funding support from NSF CAREER CCF-2143904, NSF CCF-2212066, NSF CCF-2212326, NSF IIS 2402950, ONR N000142512339, DARPA HR0011578254, and Google Research Scholar Award. We also thank Shuo Xie (TTIC; for fruitful discussions on optimization bias), SooMin Kwon, Huijie Zhang, Alec Xu (UMich; for help with problem formulation), Zhihui Zhu (OSU; for guidance on presentations), Zaicun Li (UMD; for mathematical insights), and Zhenyun Zhu (UMich; for scientific discussions).
@@ Appendix A Additional Experiments
Our theorem shows that even when the encoder and decoder are parametrized independently, training drives them to a symmetric (tied) solution. We confirm this empirically in Figure˜13, consistent with prior observations [kunin2019loss]. Accordingly, for Figures˜4 and 5 in the main text we train weight-tied ReLU DAEs.
@@ A.2  Denoising and Representations of Test Samples with ReLU DAE
As in [kadkhodaie2023generalization]), the ability to denoise an unseen test image is an equivalent check for generalization or overfitting, as shown in Figure˜14. Memorizing DAE (Corollary˜3.3) perfectly denoises a training image. On a test image, it still produces a training-data like output (visually “clear” but discarding input-specific information, producing high test MSE). Generalizing DAE (Corollary˜3.3) denoises both while preserving input-specific structure.
@@ A.3  Connection between ReLU DAE and Real-World Models
In this section, we demonstrate that our ReLU model exhibits piecewise linearity, consistent with observations in real-world models [lukoianov2025locality]. Consequently, it can be viewed as a localized approximation of these counterparts: a large model can implement the mechanisms of Corollary˜3.2 and Corollary˜3.3 in distinct local regions [ross2024geometric], thereby simultaneously generalizing and memorizing. We verify this via SVD analysis of the Jacobian [kadkhodaie2023generalization, achilli2024losing] for SD1.4, EDM, and our ReLU DAE:
@@
  • Around generalized samples, the Jacobian matrix reflects the data structures described in corollary˜3.3. Accordingly, the model produces a smoothed result, having learned a ground-truth denoiser that incorporates the constraints of the underlying distribution [niedoba2024towards].
@@ A.4  Duplication of Training Data induces Memorization
Large-scale diffusion datasets often contain duplicates due to imperfect deduplication or aggregation from heterogeneous sources [carlini2023extracting, shi2025closer]. Such duplicates are disproportionately memorized by generative models [somepalli2023understanding, chen2024towards]. Interpolating Corollary˜3.2 and Corollary˜3.3 suggests that, when a subset is duplicated, the model tends to memorize those duplicated samples while still generalizing on the rest. We observe this behavior empirically in Figure˜17 for EDM trained on CIFAR10 with a duplicated subset (and similarly for DiT on ImageNet as in Figure˜6 ).
@@ B.1  Training and Sampling Setup for ReLU DAEs
Sampling. We train a set of DAEs with VE noise scheduling [song2020score] over $\sigma\in[0.02,\,2]$ and run DDIM sampling (Eq. 1).
@@ B.2  Memorization Detection Details
Collecting mem./gen. sets. For LAION-Stable Diffusion, we follow [wen2024detecting] and use publicly available prompts curated to elicit either memorization or generalization [webster2023duplication]. For CIFAR10-EDM and ImageNet-DiT, we compute the SSCD similarity [zhang2024emergence] between each generated image and its nearest neighbor in the training set; samples with similarity >0.9>\!0.9 are labeled memorized and those with similarity <0.5<\!0.5 as generalized.
Feature extraction. For EDM we extract activations at 8x8_block3.norm0; for Stable Diffusion v1.4 at up_blocks.0.resnets.2.nonlinearity; and for DiT-L/4 we use the SiLU activation in block 12 (of 24). We apply global max pooling (spatial for Stable Diffusion v1.4 and EDM; token-wise for DiT) to obtain compact representations, though detection also works even if not. Unless otherwise noted, representations are taken at DDPM timestep $t=50$, corresponding to an equivalent noise level $\sigma_{t}\approx 0.17$ [ho2020denoising].
Input: generated image $\mathbf{x}_{0}$, timestep $t$, threshold THRES
Output: intermediate representation $\mathbf{h}$, detection flag $\mathbb{I}_{\text{mem}}$
$\mathbf{x}_{t}\leftarrow\textsc{AddForwardNoise}(\mathbf{x}_{0},t)$;
$\bm{h}\leftarrow\bm{h}_{\theta}(\mathbf{x}_{t},t,\text{condition}=\varnothing)$;
 where $\bm{f}_{\theta}(\mathbf{x}_{t},t)=\bm{g}_{\theta}\!\left[\bm{h}_{\theta}(\mathbf{x}_{t},t,\varnothing)\right]$ with $\bm{g}$ and $\bm{h}$ the decoder/encoder components;
$\mathbb{I}_{\text{mem}}\leftarrow\big(\textsc{Std}(\mathbf{h})>\textsc{THRES}\big)$;
return $\mathbf{h}$, $\mathbb{I}_{\text{mem}}$;
@@
The detection metric need not be limited to standard deviation; other effective choices include the $\ell_{4}/\ell_{2}$ ratio [Vershynin2018HDP], entropy, and max-min of the representations. We found these alternatives yield similar separability between memorized and generalized samples.
@@ Appendix C Deferred Proofs
Zeroing the off-diagonal entries of $\tilde{\bm{A}}$ weakly decreases the quadratic terms and does not increase the nuclear norm (pinching [bhatia2013matrix]). Thus a minimizer can be chosen diagonal in the $\bm{U}$-basis: $\bm{A}=\bm{U}\,\mathrm{diag}(\alpha_{1},\dots,\alpha_{d})\,\bm{U}^{\top}$.
@@
(Closeness to an actual minimizer). On $\mathcal{B}_{\delta}$, if we fix the ReLU masks to be those induced by $(\bm{W}_{2}^{\star},\bm{W}_{1}^{\star})$ at the center (see Steps (1)–(2) below), then the map
@@
We simplify $\|\bm{b}\|_{F}$ by splitting out the noise. Denote $\bm{A}_{k}:=\bm{W}_{2,(k)}\bm{W}_{1,(k)}^{\top}$ and $\bm{A}:=\mathrm{blkdiag}(\bm{A}_{1},\ldots,\bm{A}_{M})$, so
@@ C.4  Proof of Smoothness with Respect to the Infinity Norm
(3) Minimality under orthogonal re-mixing. Let an equivalent optimum be obtained by block-orthogonal mixing that preserves masks. Each active block is conjugated to $\bm{Q}_{i}^{\top}\!\big(a_{i}\,\bm{x}_{i}\bm{x}_{i}^{\top}+\lambda\bm{I}\big)\bm{Q}_{i}.$ While eigenvalues are invariant, $(1,1)$ (hence $\infty\to\infty$) norms are sensitive to densification. The memorized alignment keeps the block rank-1 plus diagonal along $\bm{x}_{i}$, which minimizes absolute column/row sums; mixing spreads mass across coordinates and (weakly) increases the max column/row sum. Therefore the memorized choice minimizes $L_{\infty}$ among all such equivalents [xie2024implicit]. ∎

### v2

@@ 3.3  Case 3: Hybrid of Memorization and Generalization with Imbalanced Data
Large-scale diffusion datasets often contain duplicates due to imperfect curation or heterogeneous aggregation [13]; such samples are more easily memorized [73] (See Section˜A.4 for more discussions). We model this by allowing duplicated (rank-1) clusters alongside well-sampled, nondegenerate clusters, so the DAE can admit local minimizers that mix memorization and generalization blocks:
@@ 4.1  Representation-Based Memorization Detection
Building on our theoretical insights, we investigate whether memorization can be detected directly from internal representations. Prior work has largely focused on how certain prompts trigger memorization and often relies on those for detection [94, 38, 65]. Representative approaches include: (i) Density-based: detecting samples that are generated disproportionately frequently under a prompt [13]; and (ii) Norm-based: comparing conditional vs. unconditional scores [94] and (iii) Attention-based: locating anomaly in the cross-attention induced by memorized prompts [35, 14]. A notable exception is (iv) a landscape-based method of [67], which evaluates memorization using local score-function geometry around a generated sample. Their method makes detection prompt-free, but is still based on output space.
@@
[13] 0.498 0.020 3.724 N/A N/A
[94] 0.986 0.961 0.134 N/A N/A
[35] 0.957 0.500 0.009 N/A N/A
[67] $\checkmark$ 0.956 0.915 0.545 0.971 0.528 0.031 0.713 0.013 0.071
@@ 4.2  Representation-Space Steering for Interpretable Image Editing
We note that we do not intend to compete with existing outstanding steering methods, as several such approaches have already demonstrated impressive empirical success [34, 111, 27, 40, 17]. Rather, our focus is on showing how steering reveals the dual relationship between representation structure and generation behavior. In particular, when the model generalizes, its representations form compositional and interpretable spaces, enabling continuous and controllable edits.
@@ 5.1  Analysis of Learning Diffusion Models with Specific Model Parameterizations
There has been a large body of work analyzing the learning of diffusion models [91, 16, 55, 100]. Recently, more attention has turned to when and how they overfit or generalize: [50, 11] use random-feature assumptions, while [89, 12] studied empirical denoisers with learnable attractors; [95, 19, 102] investigated smoothing effects induced by learning rates and weight decay that promote generalization. These works are theoretically rigorous but often lack real-world validation.
@@ 5.2  Memorization and Generalization with Analytical DMs
Recent works characterize how architectural or inductive biases can push empirical scores toward more generalizable solutions. For instance, [68, 57] constructed closed-form diffusion models from data; [59, 60, 41, 90] imposed locality or translation-equivariance constraints to mimic U-Net behavior; and [39, 3, 25] analyzed architectural biases of CNNs and DiTs. [5] empirically evaluated the impact of various regularization schemes.
@@
[64, 2, 63] model imperfect training and sampling jointly as an AM recall process, viewing novel image generation as new attraction basins and memorization as perfect recalls [10, 58]. However, this perspective can understate the role of learned neural networks in enabling generalization.
@@ 5.3  Studies on Representation Learning of Diffusion Models
Concurrent work studies co-emerging representation learning [46, 33, 101] with distribution learning in diffusion models. As recent works [18, 97] re-emphasize that the diffusion objective is fundamentally a self-supervised autoencoder loss [84, 85, 7], which induces encoder-decoder behavior [18] and the model autonomously learns informative features for downstream tasks [6, 96, 112]. Moreover, supervising the representations can accelerate training [104, 88, 71], and different representation behaviors correlate with different degrees of overfitting [54].
@@ 6  Discussion
Our Final Thoughts Diffusion models generalize mainly because, under the self-supervised denoising objective, neural networks are driven to learn and exploit the underlying (low-dimensional) structures of the data distribution. This capability is reflected internally through the emergence of semantic low-dimensional representation spaces (or auto-encoding [7]): the network effectively processes/projects noisy inputs with respect to learned structures, which underlie its compressing and denoising behavior [51, 39, 82]. In this sense, learning a good (balanced and semantic) representation is a useful indicator of model generalizability. For our theoretical analysis, we focused on a simplified setting designed for analytical tractability, which nonetheless yields foundational intuition into how real-world models function. Specifically, by examining a two-layer ReLU DAE trained on data drawn from a separable MoG distribution: a setting where denoising, score learning, and representation learning are all well defined. We demonstrate that the model learns to map inputs from the same Gaussian mode to the same ReLU mask (i.e., the same subset of active neurons). This selectivity [4, 75] leads a simple, interpretable form of representation learning, and we view it as a fundamental reason why neural networks effectively adapt to structured data distributions.
@@ Acknowledgments
We acknowledge funding support from NSF CAREER CCF-2143904, NSF CCF-2212066, NSF CCF-2212326, NSF IIS 2402950, ONR N000142512339, DARPA HR0011578254, and Google Research Scholar Award. We also thank Shuo Xie (TTIC; for fruitful discussions on optimization bias), SooMin Kwon, Huijie Zhang, Alec Xu, Prof. Laura Balzano (UMich; for help with problem formulation), Prof. Zhihui Zhu (OSU; for guidance on presentations), Zaicun Li (UMD; for mathematical insights), and Zhenyun Zhu (UMich; for scientific discussions).
@@ Appendix A Additional Experiments
Our theorem shows that even when the encoder and decoder are parametrized independently, training drives them to a symmetric (tied) solution. We confirm this empirically in Figure˜13, consistent with prior observations [45]. Accordingly, for Figures˜4 and 5 in the main text we train weight-tied ReLU DAEs.
@@ A.2  Denoising and Representations of Test Samples with ReLU DAE
As in [39]), the ability to denoise an unseen test image is an equivalent check for generalization or overfitting, as shown in Figure˜14. Memorizing DAE (Corollary˜3.3) perfectly denoises a training image. On a test image, it still produces a training-data like output (visually “clear” but discarding input-specific information, producing high test MSE). Generalizing DAE (Corollary˜3.3) denoises both while preserving input-specific structure.
@@ A.3  Connection between ReLU DAE and Real-World Models
In this section, we demonstrate that our ReLU model exhibits piecewise linearity, consistent with observations in real-world models [57]. Consequently, it can be viewed as a localized approximation of these counterparts: a large model can implement the mechanisms of Corollary˜3.2 and Corollary˜3.3 in distinct local regions [67], thereby simultaneously generalizing and memorizing. We verify this via SVD analysis of the Jacobian [39, 1] for SD1.4, EDM, and our ReLU DAE:
@@
  • Around generalized samples, the Jacobian matrix reflects the data structures described in corollary˜3.3. Accordingly, the model produces a smoothed result, having learned a ground-truth denoiser that incorporates the constraints of the underlying distribution [60].
@@ A.4  Duplication of Training Data induces Memorization
Large-scale diffusion datasets often contain duplicates due to imperfect deduplication or aggregation from heterogeneous sources [13, 70]. Such duplicates are disproportionately memorized by generative models [73, 15]. Interpolating Corollary˜3.2 and Corollary˜3.3 suggests that, when a subset is duplicated, the model tends to memorize those duplicated samples while still generalizing on the rest. We observe this behavior empirically in Figure˜17 for EDM trained on CIFAR10 with a duplicated subset (and similarly for DiT on ImageNet as in Figure˜6 ).
@@ B.1  Training and Sampling Setup for ReLU DAEs
Sampling. We train a set of DAEs with VE noise scheduling [76] over $\sigma\in[0.02,\,2]$ and run DDIM sampling (Eq. 1).
@@ B.2  Memorization Detection Details
Collecting mem./gen. sets. For LAION-Stable Diffusion, we follow [94] and use publicly available prompts curated to elicit either memorization or generalization [92]. For CIFAR10-EDM and ImageNet-DiT, we compute the SSCD similarity [110] between each generated image and its nearest neighbor in the training set; samples with similarity >0.9>\!0.9 are labeled memorized and those with similarity <0.5<\!0.5 as generalized.
Feature extraction. For EDM we extract activations at 8x8_block3.norm0; for Stable Diffusion v1.4 at up_blocks.0.resnets.2.nonlinearity; and for DiT-L/4 we use the SiLU activation in block 12 (of 24). We apply global max pooling (spatial for Stable Diffusion v1.4 and EDM; token-wise for DiT) to obtain compact representations, though detection also works even if not. Unless otherwise noted, representations are taken at DDPM timestep $t=50$, corresponding to an equivalent noise level $\sigma_{t}\approx 0.17$ [36].
Input: generated image $\bm{x}_{0}$, timestep $t$, threshold THRES
Output: intermediate representation $\bm{h}$, detection flag $\mathbb{I}_{\text{mem}}$
$\bm{x}_{t}\leftarrow\textsc{AddForwardNoise}(\bm{x}_{0},t)$;
$\bm{h}\leftarrow\bm{h}_{\theta}(\bm{x}_{t},t,\text{condition}=\varnothing)$;
 where $\bm{f}_{\theta}(\bm{x}_{t},t)=\bm{g}_{\theta}\!\left[\bm{h}_{\theta}(\bm{x}_{t},t,\varnothing)\right]$ with $\bm{g}$ and $\bm{h}$ the decoder/encoder components;
$\mathbb{I}_{\text{mem}}\leftarrow\big(\textsc{Std}(\bm{h})>\textsc{THRES}\big)$;
return $\bm{h}$, $\mathbb{I}_{\text{mem}}$;
@@
The detection metric need not be limited to standard deviation; other effective choices include the $\ell_{4}/\ell_{2}$ ratio [83], entropy, and max-min of the representations. We found these alternatives yield similar separability between memorized and generalized samples.
@@ Appendix C Deferred Proofs
Zeroing the off-diagonal entries of $\tilde{\bm{A}}$ weakly decreases the quadratic terms and does not increase the nuclear norm (pinching [9]). Thus a minimizer can be chosen diagonal in the $\bm{U}$-basis: $\bm{A}=\bm{U}\,\mathrm{diag}(\alpha_{1},\dots,\alpha_{d})\,\bm{U}^{\top}$.
@@
Furthermore, the constructed minimizer is close to an actual minimizer: On $\mathcal{B}_{\delta}$, if we fix the ReLU masks to be those induced by $(\bm{W}_{2}^{\star},\bm{W}_{1}^{\star})$ at the center (see Steps (1)–(2) below), then the map
@@
We further simplify $\|\bm{b}\|_{F}$ by splitting out the noise. Denote $\bm{A}_{k}:=\bm{W}_{2,(k)}\bm{W}_{1,(k)}^{\top}$ and $\bm{A}:=\mathrm{blkdiag}(\bm{A}_{1},\ldots,\bm{A}_{M})$, so
@@ C.4  Proof of Smoothness with Respect to the Infinity Norm
(3) Minimality under orthogonal re-mixing. Let an equivalent optimum be obtained by block-orthogonal mixing that preserves masks. Each active block is conjugated to $\bm{Q}_{i}^{\top}\!\big(a_{i}\,\bm{x}_{i}\bm{x}_{i}^{\top}+\lambda\bm{I}\big)\bm{Q}_{i}.$ While eigenvalues are invariant, $(1,1)$ (hence $\infty\to\infty$) norms are sensitive to densification. The memorized alignment keeps the block rank-1 plus diagonal along $\bm{x}_{i}$, which minimizes absolute column/row sums; mixing spreads mass across coordinates and (weakly) increases the max column/row sum. Therefore the memorized choice minimizes $L_{\infty}$ among all such equivalents [98]. ∎